# Logical Reasoning with Relation Network for Inductive Knowledge Graph Completion

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Knowledge graphs (KGs) are a collection of factual information to describe a certain field of the real world (Bollacker-etal07Freebase, ; Lehmann-etal15DBpedia, ; Mahdisoltani-etal15YAGO, ; li2022kcube, ; li2022constructing, ; dong2023hierarchy, ), which are often formulated as $\{(h,r,t)\}$ indicating there is a relation $r$ between the head entity $h$ and tail entity $t$. Since KGs are usually incomplete (zhang2022contrastive, ; dong2023active, ), as well as the expensive costs to collect facts manually, how to automatically perform knowledge graph completion (KGC) has been a widely-studied topic in the KG community (chen2020review, ; zhang2023integrating, ; ji2021survey, ). Currently, embedding-based methods (bordes2013translating, ; yang2014embedding, ; lin2015learning, ; trouillon2016complex, ; kazemi2018simple, ) play a dominant role in the KGC problem. In general, the typical methods embed the entities and relations into a low-dimensional space with various objectives and then predict the relation between any two entities, i.e., $p(r|h,t)$. However, most embedding-based methods are merely applicable under transductive settings, where KGC models predict the missing relations among entities that are already present in the training set. When some unseen entities are newly added during testing, they have to retrain the whole KG embeddings for accurate inference, which is not feasible in practice due to the high update frequency and large sizes of many real-world KGs.
@@
Recently, there has been increasing interest on KGC under inductive setting (wang2019logic, ; wang2021relational, ), where KGC models are used to infer the relations for newly-coming entities that have never appeared in the training set. The inductive KGC is more reflective of real-world scenarios, as KGs always keep introducing new information, such as new users and products in e-commerce KGs and new molecules in biomedical KGs. Some efforts have been made to obtain the inductive embeddings for new entities using external resources, such as entity attributes (hao2020inductive, ; qin2020generative, ), textual descriptions (shi2018open, ; tang2019knowledge, ) and ontological schema (geng2021ontozsl, ). But these supplemental sources are often prohibitive to acquire, impeding their success in practice. An alternative solution is to induce entity-independent rules from the KG in either statistical (galarraga2013amie, ; meilicke2018fine, ) or differentiable manners (yang2014embedding, ; yang2017differentiable, ; sadeghian2019drum, ). It regards inductive KGC as a rule mining problem, i.e. predicting the relationship $r$ for any entity pair $(h,t)$ based on predefined or adaptively learned rules. However, these methods often suffer from scalability issue and lack the capability of learning expressive representations for real-world datasets since different KGs have distinct rules with domain knowledge involved.
@@
As the prosperity of graph neural networks (GNNs) and their strong inductive bias (zhou2023interest, ; chen2022neighbor, ; chen2020label, ; chen2020graph, ), many works extend message-passing (MP) strategy used in GNNs to multi-relational graphs for inductive KGC (schlichtkrull2018modeling, ; teru2020inductive, ; wang2021relational, ; wang2022exploring, ). GraIL (teru2020inductive, ) is one of the pioneer work, which implicitly learns logical rules by reasoning over enclosing subgraphs. Specifically, GraIL first extracts an enclosing subgraph surrounding the target triple, and then generate the representations of entities by aggregating the structural information of its neighbors in the extracted subgraph. GraIL and its follow-up methods (chen2021topology, ; liu2021indigo, ; ontologyschema, ; lee2023ingram, ; jininductive, ; cui2022inductive, ; zhang2018link, ; zhang2020revisiting, ) have shown promising results using message passing over subgraphs for inductive KGC. However, the inductive capability of these methods is usually limited by two key issues. $(i)$ Entity representations generated by gathering the local information within the subgraph are over coarse-grained for accurate KG reasoning. It is because KGC always suffers from data sparsity, and the situation is even more difficult in inductive KGC, where the new emerging entities usually have few or even no connections to the original KG. $(ii)$ Existing message passing methods assume that messages are associated with nodes (entities) and are passed from nodes to nodes iteratively (xu2018powerful, ). Such entity-oriented message passing strategy weakens the role of relation semantics (wang2021relational, ), which violates the nature of inductive KGC since the learned logical evidence should be entity-independent.
@@
In real-world scenarios, the enclosing subgraph of a particular triple in the KG contains the logical evidence needed to deduce the relation between the target nodes (Teru2020InductiveRP, ). Taking Figure 1 $(i)$ as an example, learning the relation pattern among Elon Musk, TESLA, and California State can help the KGC model make a confident prediction of (Martin Eberhard, :Livein, California State) since “:LiveIn $\simeq$ :WorkAt $\land$ :LocatedIn”. To this end, we propose a novel message passing neural framework, i.e., NORAN, which aims to mine relation semantics for inductive KG completion. Specifically, we first provide a deep insight of formulating the problem to bridge the gap between embedding-based paradigm and inductive setting. Inspired by the findings, we propose relation network, a novel relation-induced graph derived from the original KG. The reconstructed relation network provides a hyper view towards KG modeling by centering relations. More concretely, its topological structure clearly reflects the distribution of relations, i.e., each node in the relation network represents a relational instance and the links among the nodes reflect the relation correlation. Thus, modeling the relation network via message-passing networks can naturally capture relation patterns as entity-independent context information to conduct inductive KGC. In this paper, we formally define such context information as logic evidence and reveal its correlations with our relation network. Our main contribution is summarized as follows:
@@
  • We propose a novel message-passing framework, i.e., NORAN, which aims to mine latent relation semantics for inductive KG completion.
  • Extensive experiments on five real-world KG benchmarks demonstrate the superiority of our proposed NORAN over state-of-the-art competitors.
2. Preliminaries
2.1. Transductive KG Completion
2.2. Inductive KG Completion

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671911isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data mining
@@ 1. Introduction
Knowledge graphs (KGs) are a collection of factual information to describe a certain field of the real world (Bollacker-etal07Freebase, ; Lehmann-etal15DBpedia, ; Mahdisoltani-etal15YAGO, ; li2022kcube, ; li2022constructing, ; dong2023hierarchy, ), which are often formulated as $(h,r,t)$ indicating there is a relation $r$ between the head entity $h$ and tail entity $t$. Since KGs are usually incomplete (zhang2022contrastive, ; dong2023active, ), as well as the expensive costs to collect facts manually, how to automatically perform knowledge graph completion (KGC) has been a widely-studied topic in the KG community (chen2020review, ; zhang2023integrating, ; ji2021survey, ). Currently, embedding-based methods (bordes2013translating, ; yang2014embedding, ; lin2015learning, ; trouillon2016complex, ; kazemi2018simple, ) play a dominant role in KGC. The typical methods embed the entities and relations into low-dimensional spaces with various objectives and then predict the relation between any two entities. However, most embedding-based methods are merely applicable under transductive settings, where KGC models predict the missing relations among entities that are already present in the training set. When some unseen entities are newly added during testing, they have to retrain the whole KG embeddings for accurate inference, which is not feasible in practice due to the high update frequency and large sizes of real-world KGs.
@@
Recently, there has been increasing interest in KGC under inductive setting (wang2019logic, ; galkin2021nodepiece, ; wang2021relational, ), where KGC models are used to infer the relations for newly-coming entities that have never appeared in the training set. The inductive KGC is more reflective of real-world scenarios, as KGs always keep introducing new information, such as new users and products in e-commerce KGs and new molecules in biomedical KGs. Some efforts have been made to obtain the inductive embeddings for new entities using external resources, such as entity attributes (hao2020inductive, ; qin2020generative, ), textual descriptions (shi2018open, ; tang2019knowledge, ) and ontological schema (geng2021ontozsl, ). However, these supplemental sources are often prohibitive to acquire, impeding their success in practice. An alternative solution is to induce entity-independent rules from the KG in either statistical (galarraga2013amie, ; meilicke2018fine, ) or differentiable manners (yang2014embedding, ; yang2017differentiable, ; sadeghian2019drum, ). It regards inductive KG completion as a rule mining problem, i.e. predicting the relationship $r$ for any entity pair $(h,t)$ based on predefined or adaptively learned rules. However, these methods often suffer from scalability and generalizability issues since different KGs have distinct rules with domain knowledge involved.
@@
As the prosperity of graph neural networks (GNNs) and their strong inductive bias (zhou2023interest, ; chen2022neighbor, ; chen2020label, ; zhou2021temporal, ; chen2020graph, ), many works extend message-passing (MP) strategy used in GNNs to multi-relational graphs for inductive KGC (schlichtkrull2018modeling, ; teru2020inductive, ; wang2021relational, ; wang2022exploring, ). GraIL (teru2020inductive, ) is one of the pioneer work, which implicitly learns logical rules by reasoning over enclosing subgraphs. Specifically, GraIL first extracts an enclosing subgraph surrounding the target triple, and then generate the representations of entities by aggregating the structural information of its neighbors in the extracted subgraph. GraIL and its follow-up methods (chen2021topology, ; liu2021indigo, ; ontologyschema, ; lee2023ingram, ; jininductive, ; cui2022inductive, ; zhang2018link, ; zhang2020revisiting, ) have shown promising results using message passing over subgraphs for inductive KGC. However, the inductive capability of these methods is usually limited by two key issues. $(i)$ KGC always suffers from data sparsity, and the situation is even more difficult in inductive KGC, where the new emerging entities usually have few or even no connections to the original KG. $(ii)$ Existing message passing methods assume that messages are associated with entities and are passed from nodes to nodes iteratively (xu2018powerful, ). Such entity-oriented message-passing strategy weakens the role of relation semantics (wang2021relational, ), which violates the nature of inductive KGC since the learned logical evidence should be entity-independent.
@@
In real-world scenarios, the enclosing subgraph of a particular triple in the KG contains the logical evidence needed to deduce the relation between the target nodes (Teru2020InductiveRP, ). Taking Figure 1 $(i)$ as an example, learning the relation pattern among Elon Musk, TESLA, and California State can help the KGC model make a confident prediction of (Martin Eberhard, :Livein, California State) since “:LiveIn $\simeq$ :WorkAt $\land$ :LocatedIn”. To this end, we propose a novel message passing neural framework, i.e., NORAN, which aims to mine relation semantics for inductive KG completion. Specifically, we first provide a deep insight of formulating the problem to bridge the gap between embedding-based paradigm and inductive setting. Inspired by the findings, we propose relation network, a novel relation-induced graph derived from the original KG. The reconstructed relation network provides a hyper view towards KG modeling by centering relations. Its topological structure clearly reflects the distribution of relations, i.e., each node in the relation network represents a relational instance and the links among the nodes reflect the relation correlation. Thus, modeling the relation network via message-passing networks can naturally capture relation patterns as entity-independent context information to conduct inductive KGC. In this paper, we formally define such context information as logic evidence and reveal its correlations with our relation network. Our main contribution is summarized as follows:
  • We propose a novel framework, i.e., NORAN, which aims to mine latent relation semantics for inductive KG completion.
  • Extensive experiments on five real-world KG benchmarks demonstrate the superiority of our proposed NORAN.
2. Preliminaries
2.1. Transductive KG Completion
2.2. Inductive KG Completion
@@ 5.1.3. Implimentation details
All baseline methods used in this paper are based on their open-source implementations. All models are implemented in Pytorch and trained on an RTX 3090 with 24 RAM. To make a fair comparison, for all models, the embedding size of entities and relations is set as $100$ for the input, latent, and output layer. Exceptionally, PATHCON, which adopted the node labeling technique to generate the node feature matrix, has a distinctive size of input embedding that varies with datasets. We optimize all models with Adam optimizer, where the batch size is fixed at $256$. Specially, we run all models on the large OGBL dataset with a batch size of $32$ to avoid out-of-memory. We use the default Xavier initializer to initialize model parameters, and the initial learning rate is $0.005$. We apply a grid search for hyperparameter tuning. The learning rate is searched from $\{0.05,0.01,0.005,0.001\}$, the margin parameter used in all baselines, and our negative sampling training scheme from 00 to $1$. We use their best configuration as the default value for other distinctive hyperparameters in the baseline methods. To reduce the randomness, we use the random seed and report the average results of three runs.
@@ 6. Conclusions and Future work
7. Acknowledgements
The work described in this paper was fully supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU 25208322).